OpenStudy (lgbasallote):

Let P(n) be the proposition "If a and b are positive real numbers, then (a+b)^n >=a^n + b^n". Prove that P(1) is true.

OpenStudy (anonymous):

(a+b)^1=a+b a^1+b^1=a+b

OpenStudy (lgbasallote):

wait...what kind of proof is this??

OpenStudy (anonymous):

Principle of mathametical Induction(PMI)

OpenStudy (lgbasallote):

....what??

OpenStudy (anonymous):

PMI....................... http://en.wikipedia.org/wiki/Mathematical_induction

OpenStudy (bahrom7893):

lol just plug in P(1)

OpenStudy (lgbasallote):

yes...but what kind of proof is it?